# Investigating the Histogram Loss in Regression

### v1

@@ Abstract
It is becoming increasingly common in regression to train neural networks that model the entire distribution even if only the mean is required for prediction. This additional modeling often comes with performance gain and the reasons behind the improvement are not fully known. This paper investigates a recent approach to regression, the Histogram Loss, which involves learning the conditional distribution of the target variable by minimizing the cross-entropy between a target distribution and a flexible histogram prediction. We design theoretical and empirical analyses to determine why and when this performance gain appears, and how different components of the loss contribute to it. Our results suggest that the benefits of learning distributions in this setup come from improvements in optimization rather than learning a better representation. We then demonstrate the viability of the Histogram Loss in common deep learning applications without a need for costly hyperparameter tuning [^1] An implementation is available at https://github.com/marthawhite/Histogram_loss.
@@ 2.2 Target Distributions
$q(y)=\frac{1}{Z\sigma\sqrt{2\pi}}e^{-\frac{(y-\mu)^{2}}{2\sigma^{2}}}$
where $Z=\frac{1}{2}(\text{erf}\left(\frac{b-\mu}{\sqrt{2}\sigma}\right)-\text{erf}\left(\frac{a-\mu}{\sqrt{2}\sigma}\right))$, and HL has
$c_{i}=\tfrac{1}{2Z}\left(\text{erf}\left(\frac{l_{i}+w_{i}-\mu}{\sqrt{2}\sigma}\right)-\text{erf}\left(\frac{l_{i}-\mu}{\sqrt{2}\sigma}\right)\right).$
This distribution enables smoothing over $Y$, through the variance parameter $\sigma^{2}$. We call this loss HL-Gaussian, defined by number of bins $k$ and variance $\sigma^{2}$. Based on positive empirical performance, it will be the main HL loss that we advocate for and analyze.
@@
$\!\!\lim_{a\rightarrow 0}\!HL(\delta_{a,j},h_{\mathbf{x}_{j}})$ $=-\log h_{i_{j}}(\mathbf{x}_{j}).$
@@
$\mathop{\mathrm{argmin}}_{h_{1},\ldots,h_{k}}-\sum_{j=1}^{n}\log h_{i_{j}}(\mathbf{x}_{j})$ $=\mathop{\mathrm{argmin}}_{h_{1},\ldots,h_{k}}-\sum_{j=1}^{n}\log h_{\mathbf{x}_{j}}(y_{j}).$
Such a delta distribution on $Y|\mathbf{x}$ results in one coefficient $c_{i}$ being 1, reflecting the distributional assumption that $Y$ is certainly in a bin. In the experiments, we compare to this loss, which we call HL-OneBin.
Mixture of Dirac delta and uniform on $Y|\mathbf{x}$ (HL-Uniform).    Using a similar analysis to above, $q(y)$ can be considered as a mixture between $\delta_{a,j}(y)$ and a uniform distribution. For a weighting of $\epsilon$ on the uniform distribution, the resulting loss HL-Uniform has $c_{i}=\epsilon$ for $i\neq i_{j}$, and $c_{i_{j}}=1-(k-1)\epsilon$.
@@ 3.3 Bias of the Histogram Loss
In this setup the predictor that is Bayes optimal in squared error predicts $\operatorname{\mathbb{E}}_{p}[y]$. We define the bias of a predicted distribution $h^{*}$ as $\operatorname{\mathbb{E}}_{h^{*}}[z]-\operatorname{\mathbb{E}}_{p}[y]$. The following proposition characterizes the bias incurred from predicting a histogram. The proof is in Appendix A.
@@
    If the form of $h$ is not restricted, the minimizer of HL is $s(z)=\int p(y)q_{y}(z)\mathop{}\!\mathrm{d}y$, i.e., the result of smoothing $p$ with $q$.
@@
    The bias of $s$ is $\operatorname{\mathbb{E}}_{y\sim p}[\operatorname{\mathbb{E}}_{z\sim q_{y}}[z]-y]$ and if the target distribution is Gaussian, $s$ is unbiased.
@@
    If the form of $h$ is restricted to a histogram, the minimizer of HL is $h^{*}$ where the probability in each bin is equal to the probability of $s$ in the range of that bin.
@@
    If the target distribution is Gaussian, the bias of $h^{*}$ is at most half of the bin width and this bound is tight.
@@ 3.4 Bound on Prediction Error
The proof is in Appendix A. The two quantities in the minimum are obtained using Pinsker’s inequality and Bretagnolle–Huber inequality and yield tighter bounds for small and large values of KL-divergence respectively (Lattimore and Szepesvári, 2020; Canonne, 2022). A model trained with HL reduces $D_{KL}(q||h_{\mathbf{x}})$ for each training data point and, if the mean of the target distribution is close to the original label, the bound above shows that the mean of predicted distribution is a good predictor for the original labels in the training set if the model achieves a low training loss.
@@ 4.1 Overview of Empirical Study
The purpose of our experiments is to first show proof of concept studies and then provide insights on the role of factors that differentiate HL from the $\ell_{2}$ loss. The main goal is to understand when and why one would opt for HL. The datasets and the algorithms used in our experiments are described in Sections 4.2 and 4.3 respectively.
@@
In Section 4.5, we investigate the impact of discretization bias on both HL-Gaussian and HL-OneBin. Previously we offered a theoretical discussion on the effect of the target distribution and the variance parameter on this bias in Section 3.3. The results in 4.5 show that this is indeed an important factor at play, and can explain most of the difference between the behaviors of HL-OneBin and HL-Gaussian. We further evaluate the models with different numbers of bins and variance parameter to see if a bias-variance tradeoff is at work.
@@ 4.8 Corrupted Targets and Input Perturbations
Figure 9: (a) Corrupted targets on CT Scan. The lines show test errors. The $\ell_{1}$ loss was robust to corrupted targets, and the $\ell_{2}$ loss was affected the most. (b) Sensitivity results on CT Scan. The horizontal axis shows the sensitivity of the model’s output to input perturbations (left means less sensitive) and the vertical axis shows the test error (lower is better). HL-Gaussian showed less sensitivity and lower test errors than $\ell_{1}$ and $\ell_{2}$.
@@ 6.2 Future Work
In terms of applications, a future direction would be finding problem-specific or even adaptive discritization methods. Breaking down the target range to even bins might not be reasonable in some cases. For example, consider the task of predicting the number of views of an online video. Presence of viral videos that heavily stretch the range of targets in the dataset will face a model designer with a trilemma: (1) choosing a wide support that covers the targets and finely discretizing it which results in a network with an enormous last layer, (2) lowering the number of bins and turning to a coarse discretization which loses information about small differences between the targets, and (3) truncating the support which requires drawing a line between informative data points and outliers and may also excessively limit the range of values the network can predict. Finding a reasonable support and the best way to discretize it are left for future work.
A Proofs
@@
The second line above proves Statement 2 and the third line shows Statement 3 for Gaussian target distribution.
@@
The histogram density that minimizes the cross-entropy to $s$ is one where the probability in each bin is equal to the probability of $s$ in the range of that bin. We call this density $h^{*}$ and find the difference between the expected values of $h^{*}$ and $p$.
We consider the case where all bins have equal width $w$. The center of bin $i$ and the probability of $h^{*}$ in bin $i$ are denoted by $m_{i}$ and $h^{*}_{i}$ respectively and $\mathcal{S}_{i}\coloneqq[m_{i}-\frac{w}{2},m_{i}+\frac{w}{2})$ is the range of bin $i$.
@@
However, as long as $\delta>0$, each term in the sum is positive and the sum is positive (because it is a part of a Gaussian pdf) and, when $\delta<0$ the terms in the sum become negative (because the integral bounds are flipped). So the two terms in the sum will have opposite signs and will not add to each other, and the error is bounded by $(-\frac{w}{2},\frac{w}{2})$. This is a tight bound since the worst case is when $\sigma\rightarrow 0$ (and the loss becomes HL-OneBin) and $\delta=\pm\frac{w}{2}$. Then $h^{*}$ will have all its density on the bin that contains $y$ and the error is $m_{0}-y=-\delta=\mp\frac{w}{2}$ (Statement 4).
@@ B Bias Simulation
C Empirical Study - Details
@@ E.4 Optimization Properties
F Robustness to Corrupted Targets
@@
G Sensitivity to Input Perturbations

### v2

@@ Abstract
It is becoming increasingly common in regression to train neural networks that model the entire distribution even if only the mean is required for prediction. This additional modeling often comes with performance gain and the reasons behind the improvement are not fully known. This paper investigates a recent approach to regression, the Histogram Loss, which involves learning the conditional distribution of the target variable by minimizing the cross-entropy between a target distribution and a flexible histogram prediction. We design theoretical and empirical analyses to determine why and when this performance gain appears, and how different components of the loss contribute to it. Our results suggest that the benefits of learning distributions in this setup come from improvements in optimization rather than modelling extra information. We then demonstrate the viability of the Histogram Loss in common deep learning applications without a need for costly hyperparameter tuning [^1] An implementation is available at https://github.com/marthawhite/Histogram_loss.
@@ 2.2 Target Distributions
$q(y)=\frac{1}{Z\sigma\sqrt{2\pi}}e^{-\frac{(y-y_{j})^{2}}{2\sigma^{2}}}$
where $Z=\frac{1}{2}(\text{erf}\left(\frac{b-y_{j}}{\sqrt{2}\sigma}\right)-\text{erf}\left(\frac{a-y_{j}}{\sqrt{2}\sigma}\right))$, and HL has
$c_{i}=\tfrac{1}{2Z}\left(\text{erf}\left(\frac{l_{i}+w_{i}-y_{j}}{\sqrt{2}\sigma}\right)-\text{erf}\left(\frac{l_{i}-y_{j}}{\sqrt{2}\sigma}\right)\right).$
This distribution enables smoothing over $Y$, through the variance parameter $\sigma^{2}$. We call this loss HL-Gaussian, defined by number of bins $k$ and variance $\sigma^{2}$. Based on positive empirical performance, it will be the main HL loss that we advocate for and analyze. Combined with Equation (2), the final form for this loss is:
[…]
@@
$\!\!\lim_{a\rightarrow 0}\!HL(\delta_{a,j},h_{\mathbf{x}_{j}})$ $=-\log h_{\text{bin}_{j}}(\mathbf{x}_{j}).$
@@
$\mathop{\mathrm{argmin}}_{h_{1},\ldots,h_{k}}-\sum_{j=1}^{n}\log h_{\text{bin}_{j}}(\mathbf{x}_{j})$ $=\mathop{\mathrm{argmin}}_{h_{1},\ldots,h_{k}}-\sum_{j=1}^{n}\log h_{\mathbf{x}_{j}}(y_{j}).$
Such a delta distribution on $Y|\mathbf{x}$ results in one coefficient $c_{i}$ being 1, reflecting the distributional assumption that $Y$ is certainly in a bin. In the experiments, we compare to this loss, which we call HL-OneBin. This makes the final form of the loss:
[…]
Mixture of Dirac delta and uniform on $Y|\mathbf{x}$ (HL-Uniform).    Using a similar analysis to above, $q(y)$ can be considered as a mixture between $\delta_{a,j}(y)$ and a uniform distribution. For a weighting of $\epsilon$ on the uniform distribution, the resulting loss HL-Uniform has $c_{i}=\epsilon$ for $i\neq i_{j}$, and $c_{i_{j}}=1-(k-1)\epsilon$. The loss can be expressed as:
[…]
@@ 3.3 Bias of the Histogram Loss
In this setup the predictor that is Bayes optimal in squared error predicts $\operatorname{\mathbb{E}}_{p}[y]$. We define the bias of a predicted distribution $h^{*}$ as $\text{Bias}(h^{*},p)=\operatorname{\mathbb{E}}_{h^{*}}[z]-\operatorname{\mathbb{E}}_{p}[y]$. The following proposition characterizes the bias incurred from predicting a histogram. The proof is in Appendix A.
@@
    If the form of $h$ is not restricted, the minimizer of the KL-divergence is the result of smoothing $p$ with $q$. Formally, $\mathop{\mathrm{argmin}}_{h\in\mathcal{P}({\mathbb{R}})}\operatorname{\mathbb{E}}_{y\sim p}[D_{KL}(q_{y}||h)]=s$ where $s(z)=\int p(y)q_{y}(z)\mathop{}\!\mathrm{d}y$.
@@
    The bias of $s$ is the expected difference between the scalar target and the mean of the target distribution constructed from it. If the target distribution is Gaussian, $s$ is unbiased. Formally, $\text{Bias}(s,p)=\operatorname{\mathbb{E}}_{y\sim p}[\operatorname{\mathbb{E}}_{z\sim q_{y}}[z]-y]$.
@@
    If the form of $h$ is restricted to a histogram with bins of equal widths, the minimizer of HL is $h^{*}$ where the probability in each bin is equal to the probability of $s$ in the range of that bin. Formally, $\mathop{\mathrm{argmin}}_{h\in\mathcal{P}({\mathbb{N}})}\operatorname{\mathbb{E}}_{y\sim p}[\text{HL}(q_{y}||h)]=h^{*}$ where $h^{*}_{i}=\int_{l_{i}}^{l_{i}+w}s(z)\mathop{}\!\mathrm{d}z$.
@@
    If the target distribution is Gaussian, the bias of $h^{*}$ is at most half of the bin width and this bound is tight, i.e. $|\text{Bias}(h^{*},p)|\leq w/2$.
@@ 3.4 Bound on Prediction Error
Figure 3: The behavior of the bound in Equation (7) on an example with Dirac delta functions. LHS shows the left-hand side, and the other two curves show the bounds obtained with Pinsker’s and BH inequality.
The proof is in Appendix A. A model trained with HL reduces $D_{KL}(q||h_{\mathbf{x}})$ for each training data point and, if the mean of the target distribution is close to the original label, the bound above shows that the mean of predicted distribution is a good predictor for the original labels in the training set if the model achieves a low training loss.
The two quantities in the minimum are obtained using Pinsker’s inequality and Bretagnolle–Huber (BH) inequality and yield tighter bounds for small and large values of KL-divergence respectively (Lattimore and Szepesvári, 2020; Canonne, 2022). To better understand these terms and the overall bound, consider the example where $a=-1,b=1,q_{y}=\delta(-1),h_{\mathbf{x}}=\epsilon\delta(-1)+(1-\epsilon)\delta(1)$ for $\epsilon\in(0,1)$. As $\epsilon$ rises, the two distributions become more similar to each other and both the squared error between their means (the left-hand side in Equation (7)) and the KL-divergence between the distributions decreases. Figure 3 shows the bounds obtained with Pinsker’s inequality and BH inequality (the first and the second term in the minimum) as well as the left-hand side. Both bounds decrease to zero as the two distributions approach each other. Pinsker’s inequality is preferred when the distributions are closer to each other. For small values of $\epsilon$, the bound using Pinsker’s inequality exceeds a trivial bound of $(b-a)^{2}=4$, while the bound with BH inequality never exceeds this value.
@@ 4 An Empirical Study on the Histogram Loss
4.1 A Synthetic Experiment
We start with an experiment on sine functions to understand the difference between HL-Gaussian and the $\ell_{2}$ loss in a controlled setting. The goal is to see how the convergence rates of the two losses compare as the frequency or the offset of the target changes. The function we want to fit has an input $x\in\mathbb{R}$ and a deterministic target $\sin(\alpha x)+\beta$ where $\alpha,\beta\in\mathbb{R}$ determine the frequency and the offset respectively. We pick 500 points on a grid from the range $[-\pi,\pi]$ and generate the corresponding targets. An example is depicted as the blue dots in Figure 4 (Left). We use a neural network with two hidden layers of width 1024 and leaky ReLU activation and train it with Adam optimizer and report the MSE on the same 500 points during the training. When comparing HL-Gaussian and $\ell_{2}$, we try a wide range of step-size values in $\{10^{-5},10^{-4},\cdots,10^{-1}\}$ for each method and show the curve with the lowest area under curve during 1000 iterations.
In the first part of the experiment we fix $\beta$ to 0 and vary the frequency $\alpha\in\{1,10,20\}$. The range of the target is $[-1,1]$ and for HL-Gaussian we discretize the range $[-1.5,1.5]$ evenly to 100 bins and set $\sigma$ to twice the resulting bin width. As shown in Figure 4 (Middle), for both losses training slows down with higher target frequencies and there is a clear difference between HL-Gaussian and $\ell_{2}$ in high frequencies. We have used a sufficiently wide range of step-size values to ensure this difference cannot be eliminated with a larger step-size. The predictions obtained with HL-Gaussian and $\ell_{2}$ for $\alpha=10$ are visualized in Figure 4 (Left).
We now fix $\alpha$ to 10 and vary the offset $\beta\in\{0,1,10\}$. The range of the target is $[-1+\beta,1+\beta]$ and the support of the histogram in HL-Gaussian is $[-1.5,11.5]$ regardless of the offset. The number of bins and $\sigma$ is similar to the previous part. The learning curves are shown in Figure 4 (Right). The convergence rate of $\ell_{2}$ deteriorates as the offset rises while the behavior of HL-Gaussian is not affected.
Overall, the results in this synthetic experiment suggest that, compared to the $\ell_{2}$ loss, HL-Gaussian is especially helpful for convergence when the target has high frequency components or is far from zero.
Figure 4: (Left) A synthetic task with freqeuncy 10 and offset 0. The blue dots show the training points and the green and red curve show the predictions obtained after training with HL-Gaussian and $\ell_{2}$ respectively. (Middle) Learning curves for different frequencies in $\{1,10,20\}$ and fixed offset 00. Green and red denote HL-Gaussian and $\ell_{2}$ and brighter shades denote higher frequencies. Each curve is averaged over 5 runs with different random initializations. (Right) Learning curves for different offsets in $\{0,1,10\}$ and fixed freqency $10$. Brighter shades show higher offsets. Altogether, HL-Gaussian trains remarkably faster than $\ell_{2}$ when the target has high frequency or is far from zero.
@@ 4.2 Overview of Empirical Study
Now we turn to the empirical study on real-world data. The purpose of these experiments is to first show proof of concept studies and then provide insights on the role of factors that differentiate HL from the $\ell_{2}$ loss. The main goal is to understand when and why one would opt for HL. The datasets and the algorithms used in our experiments are described in Sections 4.3 and 4.4 respectively.
@@
In Section 4.6, we investigate the impact of discretization bias on both HL-Gaussian and HL-OneBin. Previously we offered a theoretical discussion on the effect of the target distribution and the variance parameter on this bias in Section 3.3. The results in 4.6 show that this is indeed an important factor at play, and can explain most of the difference between the behaviors of HL-OneBin and HL-Gaussian. We further evaluate the models with different numbers of bins and variance parameter to see if a bias-variance tradeoff is at work.
@@ 4.9 Corrupted Targets and Input Perturbations
Figure 11: (Left) Corrupted targets on CT Scan. The lines show test errors. The $\ell_{1}$ loss was robust to corrupted targets, and the $\ell_{2}$ loss was affected the most. (Right) Sensitivity results on CT Scan. The horizontal axis shows the sensitivity of the model’s output to input perturbations (left means less sensitive) and the vertical axis shows the test error (lower is better). HL-Gaussian showed less sensitivity and lower test errors than $\ell_{1}$ and $\ell_{2}$.
@@ 6.1 Contributions
6.2 Limitations
Our experiments focused on neural networks and we did not study linear models in paper. Previous work on both regression and Reinforcement Learning has shown the benefits of distributional models in the context of neural networks (Bellemare et al., 2017; Imani and White, 2018; Farebrother et al., 2024) and our primary goal of our empirical study was to extend the work by (Imani and White, 2018) and better understand the underlying reasons. Nevertheless, comparing the losses in the context of linear models is a subject of independent interest and is more amenable to theoretical analysis.
Applications of regression are broad and the datasets in our empirical study and applications sections only cover a small set. Generalizing our findings to other applications such as depth estimation (Vasiljevic et al., 2019) or survival prediction (Haider et al., 2020) requires extra investigation. Similarly, emergence of new architectures and foundation models (Bommasani et al., 2021) calls for more extensive studies to ask whether our observed patterns extend to the training and fine-tuning of these models.
HL with a simple heuristic for the choice of target distribution and the extra hyperparameters performed well in our experiments. It is important to identify failure cases of these hyperparameter choices or HL in general compared to the squared error loss.
One limitation of HL is requiring knowledge about the output range. We assumed this information is available and previous work on Reinforcement Learning has treated this range as a hyperparameter (Bellemare et al., 2017).
@@ 6.3 Future Work
In terms of applications, a future direction would be finding problem-specific or even adaptive discretization methods. Breaking down the target range to even bins might not be reasonable in some cases. For example, consider the task of predicting the number of views of an online video. Presence of viral videos that heavily stretch the range of targets in the dataset will face a model designer with a trilemma: (1) choosing a wide support that covers the targets and finely discretizing it which results in a network with an enormous last layer, (2) lowering the number of bins and turning to a coarse discretization which loses information about small differences between the targets, and (3) truncating the support which requires drawing a line between informative data points and outliers and may also excessively limit the range of values the network can predict. Finding a reasonable support and the best way to discretize it are left for future work.
Appendix A Proofs
@@
The second line above proves Statement 2 and the third line shows the claim for Gaussian target distribution.
@@
We consider the case where all bins have equal width $w$. The center of bin $i$ and the probability of a histogram distribution $h$ in bin $i$ are denoted by $m_{i}$ and $h_{i}$ respectively and $\mathcal{S}_{i}\coloneqq[m_{i}-\frac{w}{2},m_{i}+\frac{w}{2})$ is the range of bin $i$.
Note that the cross-entropy between $h$ and $s$ is $\sum_{i}(\int_{\mathcal{S}_{i}}s(z)\mathop{}\!\mathrm{d}z)\log(h_{i})$. Thus the histogram density that minimizes the cross-entropy to $s$ is one where the probability in each bin is equal to the probability of $s$ in the range of that bin (Statement 3). We call this density $h^{*}$ and find the difference between the expected values of $h^{*}$ and $p$.
@@
However, as long as $\delta>0$, each term in the sum is positive and the sum is positive (because it is a part of a Gaussian pdf) and, when $\delta<0$ the terms in the sum become negative (because the integral bounds are flipped). So the two terms in the sum will have opposite signs and will not add to each other, and the error is bounded by $[-\frac{w}{2},\frac{w}{2}]$. This is a tight bound since for example when $\sigma\rightarrow 0$ (and the loss becomes HL-OneBin) and $\delta=\pm\frac{w}{2}$ then $h^{*}$ will have all its density on the bin that contains $y$ and the error is $m_{0}-y=-\delta=\mp\frac{w}{2}$ (Statement 4).
@@ Appendix B Bias Simulation
Appendix C Empirical Study - Details
@@ E.4 Optimization Properties
E.5 Robustness to Corrupted Targets
@@
E.6 Sensitivity to Input Perturbations